OpenStudy (anonymous):

Proving: \[\frac{\cos \theta \cot \theta}{\cot \theta - \cos \theta} = \frac{\cos \theta + \cos \theta \sin \theta}{1 - \sin^{2}\theta}\]

OpenStudy (anonymous):

start by writing cot(theta)=cos(theta)/sin (theta)

OpenStudy (anonymous):

i get the answer of \[\frac{\cos \theta \cot \theta}{\cot \theta - \cos \theta}\] but it is inverse of equation in right side.

OpenStudy (anonymous):

take a least common factor/multiple in denominator and pull out cos (theta) as a common factor from denominator, it should give you the right side of the equation

OpenStudy (anonymous):

i have given you pretty much a direct hint, just rework some stuff : )

OpenStudy (anonymous):

post your work, i can cross check

OpenStudy (anonymous):